OpenStudy (anonymous):

Phillip bought 12 used CDs and DVDs. CDs cost $2 each and DVDs cost $3 each. He spent $31 not including tax. How many DVDs did Phillip buy? (2 points) Question 9 options: 1) 10 2) 7 3) 5 4) 2

OpenStudy (anonymous):

2C+3D=31 C+D=12 2C+2D=24 D=7, so answer 2)